Valencia1245:

Solve for x: (x+4) ---- = 2. 3 x = −2 x = 2 x = 2/3 x =-10/3

Zepdrix:

Hmm, something didn't copy/paste correctly. :o

Valencia1245:

I edited it

Zepdrix:

The start is still messed up :\ Solve for x: =2 what is on the left side of that equals?

Valencia1245:

I edited it again

Zepdrix:

\[\large\rm \frac{x+4}{3}=2\]like this?

Valencia1245:

yeah

Zepdrix:

x has a bunch of stuff being applied to it, something dividing it, something adding to it. To solve for x, we need to "undo" all these operations by applying their inverse. We'll start by undoing the divide by 3, by multiplying both sides by 3,\[\large\rm \color{royalblue}{3}\cdot\frac{x+4}{3}=2\cdot\color{royalblue}{3}\]The multiply by 3 and divide by 3 cancel out on the left side of the equation,\[\large\rm \cancel{\color{royalblue}{3}}\cdot\frac{x+4}{\cancel{3}}=2\cdot\color{royalblue}{3}\]And multiplying out the right side gives us,\[\large\rm x+4=6\]

Zepdrix:

Understand all of that? Think you can handle the next step? Undo the addition using subtraction.

Valencia1245:

little

Zepdrix:

Subtract 4 from each side,\[\large\rm x+4\color{red}{-4}=6\color{red}{-4}\]\[\large\rm x\cancel{+4}\cancel{\color{red}{-4}}=6\color{red}{-4}\]\[\large\rm x=6\color{red}{-4}\]

Valencia1245:

ok